# Age-Based Scheduling for Mobile Edge Computing: A Deep Reinforcement Learning Approach

### 01 Dec 2023 v1

@@ II Related Work
Since the first introduction of AoI in [6], many researchers have analyzed the AoI in various systems based on the methodology from queueing theory. The authors in [5] derived analytical expressions for AoI in $M/M/1$, $M/D/1$, and $D/M/1$ systems. The results were later extended to other queueing models, such as $G/G/1$ [9], $M/G/1$ [10], and $D/G/1$ [11]. Compared to the First-Come-First-Served (FCFS) principle, the authors in [12] showed that AoI is improved in Last-Come-First-Served (LCFS) queueing systems. A variant concept of AoI, named the peak AoI (PAoI), was introduced in [13], which provides information about the maximum AoI prior to the reception of each update. The exact distribution of PAoI in a $PH/PH/1/1$ queue is analyzed in [14] and the authors in [15] computed the expected PAoI for a priority queueing system under various scenarios. In [16], the effects of packet management on both AoI and PAoI are analyzed under different pre-emption policies.
Apart from the point-to-point scenario, the AoI in queueing systems with multiple sources or multiple servers is also investigated in [17, 18, 19, 20]. In particular, the authors in [17] derived a general result for the AoI in multi-source systems under different service principles. Najm and Telatar [18] studied the AoI and PAoI in a $M/G/1/1$ model where the preemptive queue is shared by multiple streams of updates. They found that one can prioritize a stream from an age point of view by simply increasing its generation rate. The authors in [19] studied the scheduling policy that minimizes AoI in a wireless network with multiple destinations and unreliable channels. Their model was extended to a multi-source case by Sun et al. [20] and two (near) optimal scheduling policies are proposed for any time-dependent, symmetric, and non-decreasing penalty function of the ages.
Although the works mentioned above have provided numerous theoretical results and control strategies, they commonly assume the update packets sent by information sources can be directly utilized without further data processing. Moreover, their model usually ignores the resource constraints during the generation and transmission of update packets (e.g. energy consumption constraints for information sources). To fill the gap, Song et al. [21] proposed a new metric called Age of Task (AoT), which incorporates the computation time in the conventional AoI concept. By jointly considering the task scheduling and computation offloading, the authors proposed a lightweight task offline algorithm to optimize the AoT under the energy consumption constraint. However, the optimization problem considered in [21] is offline, where the arrival time of update packets is known in advance. A similar metric, named Age of Processing (AoP), is also presented in [22] and the authors utilize the Lagrangian transformation framework to solve the formulated CMDP problem. Nevertheless, the authors of [22] only consider a single WD and they adopt a generate-at-will sampling policy where the generation of updates is controllable. The authors in [23] and [24] investigated the AoI in MEC systems where the collected data can be (partially) pre-processed at the sensors. Their main contribution is to derive the analytical expressions of AoI under given statistical information. In contrast, our objective is to design a universal scheduling algorithm that can be applied to a large variety of scenarios without requiring prior knowledge of the system.
@@ III System Model and Problem Formulation
In this section, we first describe the considered MEC system and the computation offloading model. Then we introduce the concept of AoI, which measures the freshness of information contained in computation-intensive tasks. At last, we formulate an online optimization problem that aims to minimize the time-average AoI under limited spectrum bandwidth and energy consumption.
@@ III-A Network Model
As shown in Fig. 1, we consider a MEC system consisting of one base station (BS) and $N$ WDs. We assume the BS is endowed with computational resources and serves as an edge server. The WDs monitor the surrounding environment and generate a computation task containing related raw data when the environment changes. The environmental information can only be obtained after data processing. We assume the tasks are generated in an incremental way, which means the processing of one task relies on the computation results of previous tasks. Therefore, the tasks on each WD must be served in the First-Come-First-Serve (FCFS) manner.
@@ III-A2 Computation offloading
where $W_{i}(t)$ is the spectrum bandwidth allocated to WD $i$ at slot $t$, $P_{i}(t)$ is the transmission power of WD $i$, $h_{i}(t)$ is the channel gain between WD $i$ and the BS, and $\sigma^{2}$ denotes the noise power at the BS. All channels follow quasi-static flat-fading, i.e. the channel gain remains constant during each time slot, but may vary at the boundaries of slots. Hence, the amount of data offloaded from WD $i$ to the BS at slot $t$ is
@@
In practice, the duration of each time slot is relatively short so the processing of each task usually spans over multiple slots. For the convenience of analysis, we require that at most one task is processed on each WD at each time slot. In other words, if a task is completed in the middle of slot $t$, we will defer the processing of the next task to slot $t+1$. This means the totally processed data at each slot should not exceed the remaining data in the head-of-line (HOL) task. Let $d^{r}_{i}(t)$ be the amount of remaining data in the HOL task of WD $i$ at slot $t$, then $d^{l}_{i}(t)$ and $d^{o}_{i}(t)$ should satisfy
@@
Notice that constraint (7) is purely for the sake of convenience in describing the system model. We will not ensure this constraint when implementing our scheduling algorithm. Additionally, to guarantee that the lifetimes of WDs are sufficiently long, their time-average energy consumption should be lower than a threshold $E^{max}_{i}$. Let $E_{i}(t)=E^{l}_{i}(t)+E^{o}_{i}(t)$ be the energy consumption of WD $i$ at slot $t$, then this requirement can be expressed by the following constraint
@@ III-B AoI for MEC-Enabled Applications
$b_{i}(t)=\begin{cases}1,&\text{$q_{i}(t)>0$ and $d^{l}_{i}(t)+d^{o}_{i}(t)=d^{r}_{i}(t)$}\\ 0,&\text{otherwise}\end{cases},$
@@ IV-A CMDP Formulation
$\min_{\pi\in\Phi}\bar{A}^{\pi}(s)\quad\mbox{s.t. }\bar{E}_{i}^{\pi}(s)\leq E^{max}_{i},\ \forall i\in\mathcal{N},s\in\mathcal{S}.$
@@ IV-B The Lagrangian Approach
$V^{*,\lambda}(s)+v^{*,\lambda}=\min_{\alpha\in\mathcal{A}}\left\{C^{\lambda}(s,\alpha)+\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,\alpha)V^{*,\lambda}(s^{\prime})\right\}.$
@@
$Q^{*,\lambda}(s,\alpha)+v^{*,\lambda}=C^{\lambda}(s,\alpha)+\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,\alpha)V^{*,\lambda}(s^{\prime}).$
@@ IV-D Post-Decision State Learning
$p(s^{\prime}|s,\alpha)=\sum_{\tilde{s}}p_{u}(s^{\prime}|\tilde{s},\alpha)p_{k}(\tilde{s}|s,\alpha).$
@@
$\tilde{V}^{*,\lambda}(\tilde{s})=\sum_{s^{\prime}\in\mathcal{S}}p_{u}(s^{\prime}|\tilde{s})V^{*,\lambda}(s^{\prime}),$
@@
${\pi}^{*,\lambda}(s)=\operatorname*{arg\,min}_{\alpha\in\mathcal{A}}\left\{C^{\lambda}(s,\alpha)+\tilde{V}^{*,\lambda}(f_{k}(s,\alpha))\right\}.$
@@
$\sum_{t=0}^{\infty}\beta(t)^{2}+\eta(t)^{2}<\infty\mbox{ and }\lim_{t\to\infty}\frac{\eta(t)}{\beta(t)}\to 0.$
@@ V-A Average-Reward DDPG
$L^{\lambda}(\theta^{Q})=\mathbb{E}\left[\left(Q^{\lambda}(s,\alpha\,|\,\theta^{Q})-y\right)^{2}\right]$
@@
$y=C^{\lambda}(s,\alpha)-{v}_{T}^{\lambda}+{Q}_{T}^{\lambda}(s^{\prime},{\pi}_{T}^{\lambda}(s^{\prime}\,|\,\theta_{T}^{\pi})\,|\,\theta_{T}^{Q}),$
@@ V-C3 Recalculate Cost During Training
During the training process of DDPG, we use past experience stored in the replay buffer to update DNNs. The underlying assumption is that these historical data follow the same distribution. However, in our problem, the value of $\lambda$ evolves continuously, leading to different costs for the same state-action pair. Hence, the assumption that the data distribution remains unchanged no longer holds. If we use costs calculated based on outdated $\lambda$ to update the current DNNs, it may result in incorrect update directions. Fortunately, our cost function is known. Therefore, during the training process, we can recalculate the cost in the experience tuple based on the current $\lambda$, thus enhancing training stability.
@@ V-C4 Neural Network Architecture
The neural network architecture used in our algorithm is presented in Fig. 4, where “FC” stands for fully connected layer. The numbers in each block represent the dimensions of the inputs and outputs. For example, since the state of each WD is a four-tuple, the dimension of the entire system’s state is $4N$. It is evident that the architecture of the actor and value network is quite similar. The only difference lies in the addition of a sigmoid activation function at the last layer of the actor network to ensure that all outputs fall within the $[0,1]$ interval. To train these two DNNs, we use two Adam optimizers with learning rate $0.001$ (for actor network) and $0.002$ (for value network), respectively.

### 23 Feb 2024 v2

@@ II Related Work
Since the first introduction of AoI in [6], many researchers have analyzed the AoI in various systems based on the methodology from queueing theory. The authors in [5] derived analytical expressions for AoI in $M/M/1$, $M/D/1$, and $D/M/1$ systems. The results were later extended to other queueing models, such as $G/G/1$ [9], $M/G/1$ [10], and $D/G/1$ [11]. Compared to the First-Come-First-Served (FCFS) principle, the authors in [12] showed that AoI is improved in Last-Come-First-Served (LCFS) queueing systems. A variant concept of AoI, named the peak AoI (PAoI), was introduced in [13], which provides information about the maximum AoI prior to the reception of each update. The exact distribution of PAoI in a $PH/PH/1/1$ queue is analyzed in [14] and the authors in [15] computed the expected PAoI for a priority queueing system under various scenarios. In [16], the effects of packet management on both AoI and PAoI are analyzed under different pre-emption policies. Apart from the point-to-point scenario, the AoI in queueing systems with multiple sources or multiple servers is also investigated in [17, 18, 19, 20]. In particular, the authors in [17] derived a general result for the AoI in multi-source systems under different service principles. Najm and Telatar [18] studied the AoI and PAoI in a $M/G/1/1$ model where the preemptive queue is shared by multiple streams of updates. They found that one can prioritize a stream from an age point of view by simply increasing its generation rate. The authors in [19] studied the scheduling policy that minimizes AoI in a wireless network with multiple destinations and unreliable channels. Their model was extended to a multi-source case by Sun et al. [20] and two (near) optimal scheduling policies are proposed for any time-dependent, symmetric, and non-decreasing penalty function of the ages.
The aforementioned studies typically assume that update packets sent by information sources can be directly utilized. However, this assumption may not hold in various MEC applications where data processing is required to extract desirable information from raw data. To fill the gap, Song et al. [21] proposed a new metric called Age of Task (AoT), which incorporates the computation time in the conventional AoI concept. By jointly considering the task scheduling and computation offloading, the authors proposed a lightweight task offline algorithm to optimize the AoT under the energy consumption constraint. A similar metric, named Age of Processing (AoP), is also presented in [22] where the authors utilize the Lagrangian transformation framework to solve the formulated CMDP problem. The concept of AoP is also employed in [23] and [24] to facilitate the selection of multi-radio access technologies and allocation of communication and computational resources. The authors in [25] and [26] investigated the AoI in MEC systems where the collected data can be (partially) pre-processed at the sensors. Considering task-specific timeliness requirements of different applications, a low-complexity scheduling policy based on Lyapunov optimization is presented in [27]. The authors also analyzed the performance gap between the proposed policy and a theoretical lower bound of the time-average AoI penalty.
Recently, there has been extensive interest in applying DRL methods for resource management in communication systems [28, 29, 30, 31, 32]. Among these, the works most closely related to ours are [33, 34, 35, 36], as they also focus on minimizing AoI in MEC networks. Specifically, in [33], the authors considered task dependencies and proposed an offloading algorithm based on an improved dueling double deep Q-network. Jiang et al. [34] established a joint optimization model for computation offloading and transmission scheduling. To address the strong coupling between these processes, the original problem is decomposed into two stages, with a DRL algorithm proposed for each stage. Chen et al. [35] investigated the joint optimization of sensing and computation for MEC-enabled IoT networks with limited computational resources. In [36], a new metric, Age of Usage Information (AoUI), is introduced to jointly capture the freshness and usability of correlated data in IoT networks. After that, the authors employed double deep Q-network (DDQN) to construct the optimal data scheduling algorithm that minimizes AoUI. It’s worth noting that the DRL algorithms used in these works are originally designed for MDPs with completely unknown system dynamics. However, in many communication systems, part of the system dynamics is known in advance, hence directly applying these DRL algorithms would result in suboptimal results. To address this issue, this paper enhances existing DRL algorithms significantly by introducing PDSs to leverage known information about system dynamics. Experimental results demonstrate that our approach substantially outperforms existing DRL algorithms.
@@ III System Model and Problem Formulation
In this section, we first describe the considered MEC system and the computation offloading model. Then we introduce the concept of AoI, which measures the freshness of information contained in computation-intensive tasks. At last, we formulate an online optimization problem that aims to minimize the time-average AoI under limited spectrum bandwidth and energy consumption. The major notations in this paper are summarized in Table I.
TABLE I: Major Notations
Notation Description
$N$ Number of WDs.
$T$ Number of time slots.
$\Delta t$ Duration of each time slot.
$\mathcal{J}_{i}$ Index set of tasks generated by WD $i$.
$\tau_{ij}$ The $j$-th task generated by WD $i$.
$t_{ij},d_{ij}$ Generation time and data size of task $\tau_{ij}$.
$\kappa$ Number of CPU cycles required to process one bit of data.
$\gamma$ Energy efficiency factor related to the chip architecture.
$f_{i}(t)$ CPU frequency of WD $i$ at slot $t$.
$P_{i}(t)$ Transmission power of WD $i$ at slot $t$.
$h_{i}(t)$ Channel gain of WD $i$ at slot $t$.
$W_{i}(t)$ Spectrum bandwidth allocated to WD $i$ at slot $t$.
$r_{i}(t)$ Transmission rate of WD $i$ at slot $t$.
$d^{l}_{i}(t)$ Amount of locally processed data of WD $i$ at slot $t$.
$d^{o}_{i}(t)$ Amount of data transmitted from WD $i$ to the BS at slot $t$.
$d^{r}_{i}(t)$ Amount of remaining data in the HOL task of WD $i$.
$E^{l}_{i}(t)$ Energy consumption due to local computing.
$E^{o}_{i}(t)$ Energy consumption due to computation offloading.
$a_{i}(t)$ AoI of WD $i$ at slot $t$.
$b_{i}(t)$ Whether the HOL task of WD $i$ is completed at slot $t$.
$g_{i}(t)$ Whether a new task is generated by WD $i$ at slot $t$.
$q_{i}(t)$ Number of queueing tasks in WD $i$ at slot $t$.
@@ III-A Network Model
As shown in Fig. 1, we consider a MEC system consisting of one base station (BS) and $N$ WDs. We assume the BS is endowed with computational resources and serves as an edge server. The WDs monitor the surrounding environment and generate a computation task containing related raw data when the environment changes. The environmental information can only be obtained after data processing. In this paper, we focus on applications that require sequential data processing, which means the processing of one task relies on the computation results of previous tasks. One illustrative example is time-series anomaly detection, where the identification of abnormal patterns relies on historical environmental information. Consequently, the tasks on each WD must be processed in a First-Come-First-Serve (FCFS) manner. Nevertheless, our approach used to design scheduling algorithms is not bound by the FCFS property and can be easily extended to other applications where the processing of tasks is mutually independent.
@@ III-A2 Computation offloading
where $W_{i}(t)$ is the spectrum bandwidth allocated to WD $i$ at slot $t$, $P_{i}(t)$ is the transmission power of WD $i$, $h_{i}(t)$ is the channel gain between WD $i$ and the BS, and $\sigma^{2}$ denotes the noise power at the BS. All channels follow quasi-static flat-fading [42, 43, 44], i.e. the channel gain remains constant during each time slot, but may vary at the boundaries of slots. Hence, the amount of data offloaded from WD $i$ to the BS at slot $t$ is
@@
In practice, the duration of each time slot is relatively short so the processing of each task usually spans over multiple slots. For the convenience of analysis, we require that at most one task is processed on each WD at each time slot. In other words, if a task is completed in the middle of slot $t$, we will defer the processing of the next task to slot $t+1$. This means the totally processed data at each slot should not exceed the remaining data in the head-of-line (HOL) task333 In our problem, HOL task refers to the task that is currently being processed.. Let $d^{r}_{i}(t)$ be the amount of remaining data in the HOL task of WD $i$ at slot $t$, then $d^{l}_{i}(t)$ and $d^{o}_{i}(t)$ should satisfy
@@
Notice that constraint (7) is purely for the sake of convenience in describing the system model. We will provide a more detailed explanation in Section IV-B. Additionally, to guarantee that the lifetimes of WDs are sufficiently long, their time-average energy consumption should be lower than a threshold $E^{max}_{i}$. Let $E_{i}(t)=E^{l}_{i}(t)+E^{o}_{i}(t)$ be the energy consumption of WD $i$ at slot $t$, then this requirement can be expressed by the following constraint
@@ III-B AoI for MEC-Enabled Applications
$b_{i}(t)=\begin{cases}1,&\text{$q_{i}(t)>0$ and $d^{l}_{i}(t)+d^{o}_{i}(t)=d^{r}_{i}(t)$}\\ 0,&\text{otherwise}\end{cases},$ (10)
@@ IV-A CMDP Formulation
$\min_{\pi\in\Phi}\bar{A}^{\pi}(s)\quad\mbox{s.t. }\bar{E}_{i}^{\pi}(s)\leq E^{max}_{i},\ \forall i\in\mathcal{N},s\in\mathcal{S}.$ (20)
@@ IV-B The Lagrangian Approach
$V^{*,\lambda}(s)+v^{*,\lambda}=\min_{\alpha\in\mathcal{A}}\left\{C^{\lambda}(s,\alpha)+\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,\alpha)V^{*,\lambda}(s^{\prime})\right\}.$ (23)
@@
$Q^{*,\lambda}(s,\alpha)+v^{*,\lambda}=C^{\lambda}(s,\alpha)+\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,\alpha)V^{*,\lambda}(s^{\prime}).$ (24)
@@
Remark: At this stage, we can elaborate on why we need constraint (7). When describing the state transition in the previous subsection, we have categorized the process into three cases based on whether a task was completed in the preceding time slot. It’s noteworthy that without constraint (7), the system may encounter scenarios where (i) multiple tasks are completed in the previous time slot, and (ii) a new task may be partially processed. This adds considerable complexity to the state transition description as it introduces numerous new cases where the update of $a_{i}(t)$ and $d^{r}_{i}(t)$ exhibits distinct patterns. Nevertheless, conducting a comprehensive analysis of the state transition is unnecessary because we cannot directly calculate the transition probabilities due to unknown distributions of $g_{i}(t)$ and $d_{ij}$. Hence, we introduced constraint (7) to simplify the analysis of system transition. However, we will relax constraint (7) in the following algorithm design to achieve more flexible scheduling policies.
@@ IV-D Post-Decision State Learning
$p(s^{\prime}|s,\alpha)=\sum_{\tilde{s}}p_{u}(s^{\prime}|\tilde{s},\alpha)p_{k}(\tilde{s}|s,\alpha).$ (29)
@@
$\tilde{V}^{*,\lambda}(\tilde{s})=\sum_{s^{\prime}\in\mathcal{S}}p_{u}(s^{\prime}|\tilde{s})V^{*,\lambda}(s^{\prime}),$ (30)
@@
${\pi}^{*,\lambda}(s)=\operatorname*{arg\,min}_{\alpha\in\mathcal{A}}\left\{C^{\lambda}(s,\alpha)+\tilde{V}^{*,\lambda}(f_{k}(s,\alpha))\right\}.$ (33)
@@
$\sum_{t=0}^{\infty}\beta(t)^{2}+\eta(t)^{2}<\infty\mbox{ and }\lim_{t\to\infty}\frac{\eta(t)}{\beta(t)}\to 0.$ (40)
@@ V-A Average-Reward DDPG
$L^{\lambda}(\theta^{Q})=\mathbb{E}\left[\left(Q^{\lambda}(s,\alpha\,|\,\theta^{Q})-y\right)^{2}\right]$ (41)
@@
$y=C^{\lambda}(s,\alpha)-{v}_{T}^{\lambda}+{Q}_{T}^{\lambda}(s^{\prime},{\pi}_{T}^{\lambda}(s^{\prime}\,|\,\theta_{T}^{\pi})\,|\,\theta_{T}^{Q}),$ (47)
@@ V-C3 Recalculate Cost During Training
During the training process of DDPG, we use past experience stored in the replay buffer to update DNNs. The underlying assumption is that these historical data follow the same distribution. However, in our problem, the value of $\lambda$ evolves continuously, leading to different costs for the same state-action pair. Hence, the assumption that the data distribution remains unchanged no longer holds. If we use costs calculated based on outdated $\lambda$ to update the current DNNs, it may result in incorrect update directions. Fortunately, our cost function is known. Therefore, during the training process, we can recalculate the cost in the experience tuple based on the current $\lambda$, thus enhancing training stability. This also means we no longer to record the observed costs in the historical experiences.
@@ V-C4 Neural Network Architecture
The neural network architecture employed in our algorithm is presented in Fig. 4. Taking the actor network as an example. The input to this network is the system’s state, which is a collection of the individual state of each WD. Given that the state of each WD is represented as a four-tuple, the dimension of the system’s state is $4N$. This system state is then fed into a fully-connected (FC) layer comprising 128 neurons, each incorporating a rectified linear unit (ReLU) activation function. The outputs from these neurons undergo normalization via a batch norm layer to enhance the training stability and convergence of DNNs. Subsequently, the normalized results are fed into yet another FC layer and batch norm layer with the same configuration. The final layer of the actor network is an FC layer with $3N$ neurons where Sigmoid activation functions are utilized to ensure that the outputs fall within the interval $[0,1]$.
The architecture of the value network is vary similar to that of the actor network. The only difference is we only use one neuron (without activation function) at the final layer. To train these two DNNs, we employ two Adam optimizers with learning rate $0.001$ (for actor network) and $0.002$ (for value network), respectively.
@@
V-C5 Computational Complexity
In this subsection, we analyze the computational complexity of the proposed DPDS algorithm. We specifically focus on the computational workload arising from the forward and backward passes of DNNs, as the computational demands of other operations, such as sampling training batches and updating target values, are deemed negligible.
The computational complexity of DNNs is typically quantified by the number of Floating Point OPerations (FLOPs). Since the FLOPs associated with activation functions and batch norm layers are relatively minor compared to those of Fully Connected (FC) layers, our analysis exclusively considers the latter. For a single forward pass, the FLOPs of an FC layer with $n_{i}$ inputs and $n_{o}$ outputs amount to $2n_{i}n_{o}$. Accordingly, the FLOPs for the actor network and the value network per forward pass are expressed as $FLOP_{a}=2\times 4N\times 128+2\times 128\times 128+2\times 128\times 3N=1792N+32768$ and $FLOP_{v}=2\times 4N\times 128+2\times 128\times 128+2\times 128=1024N+33024$, respectively. According to the results in [66], the FLOPs associated with the backward pass are approximately twice those of the forward pass.
For DRL algorithms, the computational complexity during the execution and training phases is different. When deploying DPDS in realistic systems, obtaining the scheduling decisions in each iteration necessitates only a single forward pass of the actor network. Even for a wireless network with $N=100$ WDs, the corresponding workload is around $2.1\times 10^{5}$ FLOPs. Considering that modern GPUs, such as the Nvidia H100, can deliver up to $10^{15}$ FLOPs per second [67], the computational workload due to action selection is very light.
The training process requires slightly more computational power. Specifically, when updating the value network, we first need to calculate the loss according to equation (49), necessitating a forward pass of the target actor network, the target value network, and the value network. The obtained loss is then employed to update the value network through the backward pass. Hence the total FLOPs involved in a single update of the value network is around $FLOP_{a}+FLOP_{v}+FLOP_{v}+2\times FLOP_{v}=5888N+164864$. The FLOPs for updating the actor network can be derived similarly. In total, the computational workload for the training process in a single iteration is approximately $29619212288N+$ 296192 FLOPs.
@@ VI Numerical Results
For comparison purpose, we also attempted to implement the tabular PDS learning algorithm proposed in Section IV-D. Unfortunately, the tabular PDS learning algorithm failed to learn any effective policy for a fairly long time and the outputs seem to be totally random. The main reason for this problem is the space of the PDS is prohibitively large. Recall that the PDS for an individual WD is a four-tuple: $(\tilde{d}^{r}_{i},\tilde{a}_{i},\tilde{q}_{i},\tilde{h}_{i})$. For simplicity, we assume the max values of $\tilde{a}_{i}$ and $\tilde{q}_{i}$ are $50$ and $20$, respectively. Moreover, we discretize the data size $\tilde{d}^{r}_{i}$ and channel gain $\tilde{h}_{i}$ into $50$ and $20$ different levels. Then the number of possible PDS for an individual WD is $50\times 50\times 20\times 20=10^{6}$. By utilizing the cost function decomposition technique proposed in Section IV-F of [56], the number of possible PDS of the entire system is linear (instead of exponential) with respect to the number of WDs. Therefore, the overall number of possible PDS is approximately $15\times 10^{6}=1.5\times 10^{7}$. However, this number is still too large for the tabular PDS learning algorithm to converge. The above observation highlights the limitations of tabular PDS learning when handling practical situations with large state spaces, whcih also serves as a validation of our analysis in Section IV-E.